what maximum amount of ammonia in kilograms can be synthesized from 5.22 kg of H2 and 31.5 kg of N2?
Answer #1

Balanced chemical reaction-

N​​​​​​2 + 3H2 - 2NH3

first we find the limiting reagent than we calculate the maximun yield.

Molar mass of N​​​​​​2​​​​​- 28 g = 28/1000 = .028kg

Molar mass of H​​​​​​2- 3×2 = 6 g = 6/1000 = .006 kg

0.28 kg of N​​​​​2​​​​​ for completereaction require.006 kg of H​​​​​​2

1 kg of N​​​​​​2​​​​​ for completereaction require .006/.028 kg of H​​​​​​2

31.5 kg N​​​​​​2 for complete reaction require 0.006*31.5/.028 =6.75 kg

0.006 kg of H​​​​​​2 for completereact require 0.028 kg of N​​​​​​2

1 kg kg of H2 for complete reaction require 0.028/0.006 kg of N​​​​2

5.22 kg of H​​​​​​2 .for complete reaction require 0.028*5.22/0.006 = 24.36 kg

We have H​​​​​​2 on 5.22 kg. Means H​​​​​​2 is limiting reagent.

0.006 kg of H​​​​​​2 on complete reaction gives 0.034 kg of amonia

1 kg of H​​​​​​2 on complete reaction gives 0.034/.006 kg of ammonia

5.22kg of H​​​​​​2 .on complete reaction gives 0.034*5.22/.006 = 29.58 kg